OpenStudy (tiffanymak1996):

consider the sequence : a1= sqrt 12 a2= sqrt(12+sqrt12) a3= sqrt(12+sqrt(12+sqrt12)) a4= sqrt(12+sqrt(12+sqrt(12+sqrt12))) . . . a) find a recursion function for a(n+1) (term) b) Assuming that the sequence converges, find the limit.

OpenStudy (amistre64):

well, from the pattern, it appears that a next term is equal to the sqrt of the now term+12

OpenStudy (amistre64):

\[A_{n+1}=\sqrt{A_n+12}\] right?

OpenStudy (tiffanymak1996):

yes

OpenStudy (amistre64):

rewriting the sqrt as an exponent might be helpful ... but might not :) \[A_{n+1}=(A_n+12)^{1/2}\]

OpenStudy (amistre64):

should we try to develop an explicit formula for An?

OpenStudy (tiffanymak1996):

no, i found the limit, it's 4. Thx a lot. :)

OpenStudy (amistre64):

whew!! thats good :)